OpenStudy (anonymous):

Use linear approximation to estimate sec(44)

OpenStudy (dean.shyy):

Take a look at this: http://is.gd/I43iQe

OpenStudy (anonymous):

So this is my process for solving it according to what you showed. f(x) = sec(x) x=44 a=45 L(x)=f(a) + f'(a)(x-a) f(a)= sec(45) = 2/sqrt(2) f'(x) = secxtanx f'(45) = 2/sqrt(2) L(x) = 2/sqrt(2) + [2/sqrt(2)](44-45)

OpenStudy (anonymous):

@Dean.Shyy

OpenStudy (dean.shyy):

These may also help you: http://is.gd/ayWKgK & http://is.gd/wFcbM8 & http://is.gd/4mOktp

OpenStudy (anonymous):

What am I doing wrong? Because when I followed through I always end up with 0?

OpenStudy (dean.shyy):

Maybe you can enter your values into this: http://is.gd/nQhaEY

OpenStudy (anonymous):

Okay I think I figured it out! f(x)=sec(44) f'(x)=secxtanx f'(a)=2/sqrt(2) f(a)=2/sqrt(2) L(x)=2/sqrt(2) + 2/sqrt(2)[x-pi/4] =2/sqrt(2) + 2/sqrt(2)x - pi/sqrt(2) =2/sqrt(2) + pi/sqrt(2) - pi/sqrt(2) =2/sqrt(2)=sec(22pi/90) APPROXIMATELY = sec(pi/4) I checked it with my calculator and the answers are only about .02 off, and I think that's right. What do you think @Dean.Shyy?